What is number of possible samples from the given: N = 10, n = 3.
1
Expert's answer
2021-08-23T15:50:40-0400

Number of samples without replacement: "C_N^n=C_{10}^3=\frac{10!}{(10-3)!3!}=120."

Number of samples with replacement: "N^n=10^3=1000."